OpenStudy (anonymous):

30to the second power + 18 to the second power equals ?

OpenStudy (diyadiya):

1224

OpenStudy (anonymous):

\[30^2 = 30 \cdot 30 = 900\]\[18^2 = 18 \cdot 18 \]\[= (10 + 8)(10+8) \]\[= 100 + 2(80) + 64 = 100 + 160 + 64 = 324\] \[\implies 30^2 + 18^2 = 900 + 324 = 1224\] Didn't want to bother getting out the calculator, sorry

OpenStudy (amistre64):

thats a lot of variables ...